### 12 Jan 2024 v1

# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

Abstract
The subway system as the most important transportation for daily urban commuting is a typical non-equilibrium complex system. Unlike complex systems composed of a single type of basic units, the subway system consists of 2 types of basic units with service relationship: passengers as the demander of services and trains as the supplier of services. The main challenge of the operation of subway system arises from passenger stranding at stations, which arise from the spatiotemporal mismatch of supply scale and demand scale. There is a special phenomenon of the propagation of passenger stranding (PPS) within the stations in the system by forming stranded stations clusters under the condition of shortage of supply, which significantly reduces the efficiency of the subway system. In this study, we take the Beijing subway as an example to study the nature of PPS phenomena from a view point of statistical physics, in which both an ideal model and a realistic data-driven model are considered. The simulation results suggest that both the efficiency of subway service (as measured by passenger waiting time) and the scale of stranding demonstrate phase transition like behaviors. The phase transition point can quantitatively characterize the resilience failure threshold of subway service. PPS phenomenon encompasses a multitude of coupling effects within the subway system. The eigen microstate method is used to extracting the fundamental patterns of the formation and regulation mechanism of stranded stations clusters. The research in this study provides a theoretical basis for understanding the mechanism of service efficiency in subway networks and offers a theoretical foundation for strategies to improve service, such as topological planning and train timetable optimization. Moreover, the methodology used in present work has significant implications for service systems.
Keywords First keyword  $\cdot$ Second keyword  $\cdot$ Third keyword More
1 Introduction
Among various urban transportation modes, the subway’s travel time is the most easily controlled due to its immunity from issues like road congestion and inclement weather. The carbon emissions per passenger transported and unit distance covered by electric-powered subways are significantly lower than other urban transportation modes. As a result, subways offer an optimal solution to address common urban issues such as traffic congestion, environmental pollution, and limited efficiency of public services in large cities. The spatial extension characteristics of subways also shape the structure of large cities and the characteristics of residents’ activities, known as Transit-Oriented Development (TOD)[1] . By utilizing the economy and efficient subway travel mode, residents can opt to reside in low-cost suburbs while engaging in employment in the city center, resulting in a significant spatial decoupling between job and residence centers. As the most crucial public service system in cities, it is required that the subway system should provide efficient services to meet the highly concentrated daily commuting demands. The efficiency of a subway system is primarily characterized by its ability to fulfill the transportation demands of passengers in a timely manner, as well as minimize passenger waiting time.
Because the arrival of trains as service providers is not continuous but discrete in time, and the arrival of passengers as service demanders at the station is random, there are always random events of passenger stranded at the platform, and there is no temporal correlation between these events at stations. When the passenger arrival rate at a station exceeds the efficiency of subway’s service, the events of passenger stranding start to exhibit temporal correlation. During the morning rush hours, the Beijing subway system encounters the most severe instances of stranding, with 8.547% of stations affected (Fig.1 A illustrates the number of stations with stranding phenomena over time on Monday, May 10, 2021). The stranding phenomenon can propagate from station to station along the subway line, and these stranded stations forming a spatially correlated cluster. We call this phenomenon as the propagation of stranded passengers (PPS). PPS comprises cluster formation. As an example, the stranded stations during the morning rush hours are illustrated in Fig.1 B. The presence of temporal and spatial correlations of stranding in the subway not only leads to longer commuting times for passengers, but also undermines the advantages of subway travel and depresses societal efficiency. When the duration and spatial scale of stranding reach a substantial magnitude, the system is unable to provide adequate service supply to meet the demand. The demand scale in this case exceeds the threshold of system resilience. Large-scale passenger stranding within enclosed subway stations increases the difficulty of evacuation and raises concerns regarding public safety, such as the spread of epidemics and the risk of terrorist attacks.
The critical demand scale where the subway system fails to provide services is defined as the threshold of service resilience, which reflects the ability to meet demand and serves as a crucial parameter for quantitative evaluation in decision-making processes related to investment, insurance, risk management, and emergency preparedness. The term "resilience" originated in mechanics to describe the resistance of materials to physical impacts. In the study of complex systems, the concept of resilience can be primarily categorized into three frameworks[2] . (1) System resilience (or ecological resilience) focuses on the magnitude of the change or perturbation that a system can endure without shifting to another stable state[3] . (2) Engineering resilience is defined by the recovery rate or time[4]. (3) Adaptive resilience characterizes the capacity of socioecological systems to adapt or transform in response to unfamiliar, unexpected and extreme shocks[5] . The resilience of services can be attributed to the category of system resilience. In present work, the system’s states are characterized by its responsive properties, with demand scale serving as the stimulus and the scale of stranding representing the system’s response. The threshold of resilience is defined as the demand scale at which a significant change in response behavior occurs, at which the system can no longer provide normal service beyond a certain threshold.
The subway system is a non-equilibrium complex system composed of a vast number of interacting basic units (passengers and trains), making it a typical subject of study in statistical physics[6] . Unlike the systems composed of a single type of unit, the subway system is a service system consisting of two distinct types of units interplay through service relationship: (1) passengers as the demander of services and (2) trains as the supplier of services. A demand of service from a single passenger is a pair of stations depending on time called as origin-destination (OD) pairs, and on the supply side of services, subway trains operate according to the schedules of different lines. Stranding phenomenon of passengers at stations and the propagation of stranding within the network are the result of collective behaviors of both passengers and trains, which are from the alignment of supply and demand at each station in a timely manner, as well as the spatial correlation between stations facilitated by trains. For the condition of supply scarcity, the stranded stations become interconnected to form clusters. The generation of these clusters relies on a nucleation and growth like processes[7] , which is similar with the condensation of gas molecules attracting and coalescing into a liquid state during vapor-liquid transitions, and the growth process of magnetization in ferromagnetic materials, exhibit similarities to phase transition phenomena. To reveal the mechanism underlying the PPS phenomenon, it is essential to conduct research employing statistical physics theories and methodologies.
Generally, the subway was studied by traffic fluid model[8] or complex networks[8]. The traffic fluid model is inspired by the studies of congestion in road traffic[9; 10; 11; 12; 13; 14]. The fundamental diagram of traffic flow and velocity reflects the spontaneous phase transition of congestion[15]. Subway studies based on complex networks investigate the reliability of efficient travel paths by establishing undirected network models[16; 17] or directed network models[18], considering the network connectivity characteristics of stations. Unlike traffic flow in roads and transportation problems in complex networks, congestion phenomena in subway systems arise from the spatiotemporal matching of service and demand. Insufficient service provision results in passengers being stranded. The relaxation time of stranding phenomena at individual stations is governed by the ratio of local service supply rate and the demand rate. The service supply rate is determined by the passenger capacity of trains and the arrival rate of trains (timetable), and the demand rate is determined by the arrival rate of passengers at this station. The development patterns of TOD lead to highly spatiotemporally uneven demands for subway services due to the tidal characteristics of commuting during morning and evening rush hours[19; 20]. This makes it challenging to achieve spatial and temporal matching between service supply and demand. Besides the local mismatching of demand and supply, it is worth noting that the stranding conditions at upstream stations would reduce the supply rate of downstream station as well due to the characteristics of the train providing service along the lines, as depicted in Fig.1 D. When the train operating at almost full capacity, the number of passengers getting off the train might lower than that who want to get on which gives rise to the propagation of stranding phenomena along the line. The TOD mode contributes to a significant spatial separation between job and residence centers, resulting in a large average commuting distance for passengers which determines the spatial correlation length of such stranding phenomena.
In order to determine the resilience threshold of a system, it is necessary to investigate the response behavior of subway systems under varying demand scales. However, executing controlled variable experiments within the actual infrastructure of mega-cities, particularly when aiming to surpass the resilience threshold, encounters substantial obstacles. As a typical non-equilibrium system, a data-driven dynamic model can be established, enabling the acquisition of its response patterns under different demand scales through computer simulations. In turn, the service resilience threshold can be predicted. In present work, the system’s responsive behaviors to varying demand scales are investigated in both steady-state and non-steady-state non-equilibrium conditions. The responsive behaviors are characterized by both the scale of stranding within the system and the average waiting time of passengers. The steady-state conditions are investigated under the assumption of spatial and temporal homogeneous in both demand and supply. In this condition, the supply rate is fixed and the average demand rate (Pu) at each station is varied to reveal the general nature of the PPS phenomenon arising from supply shortage. Additionally, to simulate non-steady-state conditions, we incorporate time-dependent demand and supply rates with rush hour features in both the morning and evening, obtained from real-system data. Similar to the steady-state conditions, the supply rate is fixed as well and the demand rate, Pr is varied. Here, Pr is defined by the ratio of simulated passenger arrival rate to actual passenger arrival rates.
PPS phenomenon encompasses a multitude of coupling effects within the subway system. Extracting the fundamental patterns from these coupling effects can provide valuable insights for the development of subway infrastructure, regulation of social services, and the establishment of effective methods for emergency response and incident prediction. In the field of non-equilibrium statistical physics, eigen microstate method offers a powerful method to investigate phase transitions, critical phenomena, and dynamic evolutions in complex systems, thereby providing a new avenue for studying PPS phenomena. By applying the eigen microstates method[21; 22; 23; 24; 25; 26], fundamental patterns and their evolutionary rules can be extracted from extensive system evolution data. In present work, PPS as a class of emergent behaviors in self-organized multi-body systems are studied from a statistical physics perspective, using eigen microstates method to uncover the underlying mechanisms.
Service system can be considered as a general model for many economic and social activities of human beings. In these systems the scheduling of supply is optimized to meet the spatiotemporal characteristics of demand and PPS phenomena is a common feature of these system. Furthermore, the global operational efficiency, cost-effectiveness and prevent system from collapse are essential considerations from society management point of view. Examples of service systems include transportation systems, like subway, taxis[27] , buses[28; 29] etc; elevators in skyscrapers[30; 31; 32; 33]; medical appointment systems; logistical management[34; 35]; aircraft carriers decks management[36; 37; 38]. In present work, the subway system as a reprehensive example of service system is studied and the mechanism of PPS phenomena of service system is revealed. The results and methodologies may provide valuable insights for service systems.
Figure 1: Stranding phenomena is a major problem of subway system: (A) The temporal variation in the number of stations with stranded passengers in the subway network throughout a day on May 10, 2021; (B) Spatial distribution of stations with stranded passengers in the Beijing subway network at 8:10 AM rush hour; (C) Mechanisms of passengers being stranded at subway stations. The causes for the occurrence of stranded passengers and PPS phenomena are as follows: network topology, service trains scheduling, and spatiotemporal characteristics of demand. (D) the scheme of causing of PPS phenomena. stranding occurs at stations when passengers’ demand exceeds transportation supply. Due to the interconnected nature of the line, operation in full capacity in upstream stations would occupy the supply scale of downstream stations, leading to passengers stranded at the latter.
2 Result
To investigate the phenomenon of PPS in subway systems, we have developed a dynamic model based on an actual subway network, using Beijing as a case study. Beijing, a world-class megacity with a population of 23 million permanent residents. As of July 26, 2022, the Beijing subway network composes of 17 lines with a total of 330 operational stations and 69 transfer stations, constituting a transportation system transportation system of significant scale. The job and residence centers in Beijing exhibit characteristic features of TOD, with a substantial portion of the population relying on the subway for their daily commuting demands. In present work, a model system includes 3 east-west linear lines (Lines 1, 6, and 7), 3 north-south linear lines (Lines 4, 5, and 8), and 3 circular lines (Lines 2, 10, and 13) are selected, resulting in a subway network comprising 234 stations (N=234). Among these stations, there are 36 transfer stations and 10 terminal stations (ends), capturing the representative topological structure of the Beijing subway system. These 9 earliest established lines, operating in the core area of Beijing serve as the primary main arteries in the real subway system. This model system is an appropriate representation in terms of scale, significance, and topological structure, which enables us to capture the daily operational features. In order to concentrate on the primary physical properties and alleviate computational burden, we have excluded the newly constructed lines from our study that have not yet achieved supply-demand equilibrium.
Within this subway model, a dynamic evolutionary model system to quantitatively investigate the phenomenon of PPS is put forward. This model system comprises two types of active basic units, passengers ($N_{p}=3,811,418$) and trains ($N_{t}=5,266$) with service relationship. The service process includes the operation of trains to meet the demand of passengers. Trains operate according to a predetermined schedule, ensuring punctuality. When a train arrives at a station, passengers on board make decisions to disembark or not according to their destination demand, while passengers waiting at the station Passengers decide whether to get onboard based on the crowdedness of the compartment. Through computer simulations, the dynamic trajectories of passengers and trains can be obtained (further details can be found in Supplement Meterials, A.2). To elucidate the fundamental physical characteristics of PPS, 2 simulation systems are considered: an ideal model (System I) and a realistic model (System II). In System I, it is assumed that the rates of demand and supply are independent of time and position, and the positions of OD pairs are randomly distributed throughout the network. This model system allows us to study PPS under non-equilibrium steady-state conditions. In System II, the rate of passenger demand depending time exhibiting with rush hour patterns. The positions of OD pairs capture the tidal nature of real commuting scenarios. The simulation time is measured in seconds, with the system state recorded every 5 minutes for System I and every 10 minutes for System II.
This study employs the average waiting time of passengers and the scale of stranding as the order parameters characterizing the PPS phenomena. The waiting time of passengers, $T_{w}$, will increase due to the PPS phenomena, which serves as a crucial metric for evaluating the system’s service efficiency. To quantify the PPS phenomena in the subway network, we develop here a quantitative framework based on a physical concept of cluster, which combines evolving traffic dynamics with network structure. Instead of the commonly used structural topology, only stations in the network with stranded passengers are considered functionally connected. In this way, we can characterize and understand the dynamics of PPS phenomena through cluster formation and vanishing. The scale of stranding in the system is characterized by 3 order parameters: the total number of stations within the stranding clusters, n_stranded; the size of the largest cluster, max_cluster, represented by the number of stations within it; and the total number of stranded stations in the system, tot_stranded(quantitative definition of clusters can be found in Supplement Meterials, B.1.
2.1 System I: Steady-State features
Assuming spatial and temporal homogeneity in both supply and demand, the system can reach a non-equilibrium steady state. By studying systems in this steady state, the influence of the subway network’s topological structure on the PPS phenomenon can be revealed. In System I, supply scale remains constant by establishing a consistent train departure interval (4 minutes) from the originating station. The supply-demand ratio in the system by adjusting the demand scale, $P_{u}$ (the number of passengers entering the station per minute). The specific simulation methodology can be found in the Supplement Meterials, A.2.3. After a relaxation period of finite duration, the system reaches a non-equilibrium steady state, where the average waiting time of passengers and the values of these 3 order parameters steadied at specific values (examples in Supplement Meterials, B.2.1, Fig.B.13). The time dependent number of stranded passengers can be seen in Supplement Meterials, B.2.2, Fig.B.14, Fig.B.15. System responsive behavior to the demand scale, $P_{u}$, can be characterized by the time-averaged waiting time $t_{w}$ and the 3 order parameters that describe the scale of stranding, which is shown in Fig.2. Both $t_{w}$ and 3 order parameters exhibit similar responsive behaviors with a transition point around $P_{u}$=20. When $P_{u}$ is small, there are few stations with stranded passengers on the platforms randomly, and passengers can board the trains after a short waiting time. There is no spatial-temporal correlation between stations with stranded passengers, indicating the absence of the PPS phenomenon. Therefore, all 3 order parameters that characterize stranding scale tend to zero at small $P_{u}$ values. When $P_{u}$>20, $t_{w}$ increases abruptly with $P_{u}$, indicating a significant decrease in system efficiency. Stations with stranded passengers start to correlate with each other, forming clusters. The stranding scales indicated by all these 3 order parameters rapidly increase with increasing $P_{u}$. The behavior that significant changes in its response properties beyond a certain threshold is similar to that of a continuous phase transition. The threshold value, $P_{u}$=20, can be regarded as the transition point of this phase transition. When $P_{u}$ is below this threshold, there is no correlation in the occurrence of stranding phenomena between the stations. We call this state as a normal phase. However, when $P_{u}$ exceeds this threshold, stranded stations correlate with each other, forming persistent clusters of stations with stranded passengers that are difficult to dissipate. We call this state as a stranded phase. The system exhibits an abrupt change in its response behaviors to the demand scale exceeds the transition point, and the service efficiency dramatic decreased. We can use the value of $P_{u}$ at transition point as the threshold beyond which the system can no longer provide normal service.
Figure 2: (A) Average waiting time of passengers, $t_{w}$, as a responsive behavior of the demand scale, $P_{u}$. (B)-(D) show the curves illustrating the changes in n_stranded, max_cluster, and tot_congestion as responsive behaviors of $P_{u}$. A phase transition like behavior occurs at $P_{u}=20$.
Because there is no spatiotemporal matching issue between demand and supply in steady-state conditions, the formation of stranding and taking place of PPS phenomena solely arise from topological factors. The topological factor is characterized by the network betweenness centrality of each station. We find that when system is in the stranded phase, most of the stranded stations tend to cluster around the stations with high network betweenness centrality. To depict the dependence between the occurrence of stranding at stations on the topological factor, we sort the stations in stranded phase (taking $P_{u}$=25, 30, 35, and 40 as examples) in descending order based on the number of stranded passengers and the betweenness centrality, respectively. Using the ranks of these 2 queues as the vertical and horizontal coordinates in Fig. 3. If all stations fall on the diagonal line of this graph, the stranding phenomena and topological factors are linearly correlated. As indicated by Fig. 3 when systems are in stranded phase, the stations with higher betweenness centrality exhibit relatively higher numbers of stranded passengers and those stations with lower betweenness centrality have no passenger stranded. Taking $P_{u}$=30 as an example (blue circle), there are 33 stranded stations, with the maximum value of the vertical axis representing the ranking of passenger strand being 34. Those stations with rank above 74 which have lower betweenness centrality do not experience passenger stranding.
Generally, a station with higher betweenness centrality indicates a higher level of connectivity with the network, which would reduce the occurrence of stranding. However, results in Fig. 3 show contrary to intuition. Indeed, this outcome reflects the characteristics of subway service. In systems under steady-state condition, since the passenger demands (pairs OD stations) are randomly generated in space, it has a high probability to require transfers to different lines at stations with higher betweenness centrality. Although these stations have high connectivity, passengers have to wait for the transfer train to arrive before they can continue their travel, resulting in a delay in receiving service. It should be noted that train scheduling in different lines is independent of each other, so passengers needing transfers generally cannot receive immediate service. This is equivalent to increasing the demand volume at the transfer stations acting as hubs, leading to stranding. By analyzing the steady-state conditions, the emergence stranding phenomena and formation of clusters in the network are observed, which shows a phase transition like behavior. The mechanisms behind this behavior are the network’s topology and the independence of scheduling in different lines in subway. The steady-state conditions analyzing is used to reveal the feature of spontaneous emergence of stranded cluster in subway. However, such phenomena are not severe in real subway service, as the OD pairs of actual commuting passengers are not random. Citizens choose residence and workplace to minimize the possibility of transfers and enhance travel convenience. Hereafter, to investigate the PPS phenomenon in real systems, we will establish a computational model based on Automatic Fare Collection (AFC) data.
Figure 3: Relationship between the ranking of station betweenness centrality and the ranking of stranded passenger numbers. The betweenness centrality ranking of stations is based on the descending order of the betweenness centrality values for each station.
2.2 System II: Spatio-temporal Imbalance between Supply and Demand
Large cities exhibit a large spatial separation between job and residence centers, particularly influenced by the TOD urban development model. Citizens tend to select residence and job locations along subway lines, resulting in a spatiotemporal imbalance in commuting demands with rush hours and long-distance travel patterns. To model the non-steady-state evolution of PPS, a tidal supply-demand model at a daily time scale based on the spatiotemporal characteristics of AFC data and train timetable data is constructed. The AFC data covers the entire weekdays from May 10th to May 14th, 2021, for the nine subway lines depicted in Fig.1B of the model (more details in Supplement Meterials, A.2). The AFC data consists of 3,811,418 records of passengers ($N_{r}$) and 5,293 records of trains ($N_{ori}$). The timetable data is from Beijing Rail Transit App (more details in Supplement Meterials, A.1). The spatiotemporal mismatch between supply and demand is the primary mechanism of PPS phenomena. In the simulation, the supply scale is fixed at the real supply level while varying the demand scale to investigate the effects of service shortages on PPS. The demand scale is defined as $P_{r}=N_{d}/N_{r}$, where Nd represents the total number of passengers taking subway throughout the day in the simulation, and $N_{r}$ represents the real total number of passengers. We generate Nd OD pairs of demand by sampling the OD pairs in the same time interval based on the probability $P_{r}=N_{d}/N_{r}$ from the real data throughout the weekdays (see Supplement Meterials, A.2.4 and Fig. A.12 for parameter settings and validation).
Unlike the results obtained under steady-state conditions, the computational results based on real-system data exhibit spatiotemporal heterogeneity. In spatial dimension, there are 4 significant spatially independent clusters of stranded stations (Cluster I-IV), regardless of the demand scale, whether it is low ($P_{r}=1.0$) (Supplement Meterials, B.3.3, Fig.B.17(E) or high ($P_{r}=1.6$) (Supplement Meterials, B.3.3, Fig.B.17(J)). These clusters represent distinct commuting "corridors" connecting the residence centers located in the suburb (north: Huilongguan, Tiantongyuan; south: Daxing; east: Beijing Municipal Administrative Center and North Three Counties, including: Sanhe City, Dachang Hui Autonomous County and Xianghe County; west: Shijingshan) to the job centers situated in the city center. The formation of these commuting corridors is attributed to the mutual influence of subway construction on the spatial distribution of residential and employment areas, thereby characterizing Beijing as a Transit-Oriented Development (TOD) type city. The list contains the stations (names and labels) included in each Cluster can be found in Supplement Meterials, B.3.3, Tab.B.6,B.7. In temporal dimension, instead of steady-state features, the system exhibits characteristics of rush hours. Several time points are selected to observe PPS, as depicted in Supplement Meterials, B.3.3, Fig.8. For a demand scale of $P_{r}=1.0$, at 7:10, only a few isolated stations manifest stranding phenomena, and these stations are spatially independent. By 8:00, PPS occurs, signifying the association of stations with stranded passengers and the formation of clusters. By 8:20, the size of these clusters reach their maximum. The stranding phenomena completely vanish from the network by 13:00 (Supplement Meterials, B.3.3, Fig.8(A)-(D)). When the demand scale is high, such as $P_{r}=1.6$ (Supplement Meterials, B.3.3, Fig.8(F)-(I)), with the arrival of the morning rush hour, the system demonstrates associations between stations with stranded passengers, leading to cluster formation. As the morning rush hour subsides, the scale of PPS gradually decreased. In contrast to the case with a lower demand scale, during the onset of the evening rush hour at 13:00, the clusters generated during the morning rush hour do not entirely dissipate (Supplement Meterials, B.3.3, Fig.8(I)). The system is always in the working state with high stranding phenomenon, and the efficiency of service delivery is quite low.
In the context of spatiotemporal heterogeneity driven by real-system data, both passenger waiting time and 3 order parameters fail to reach steady states, exhibiting dynamic fluctuations over time. Consequently, the method to determining the transition point of the system, as depicted in Figure 2, becomes unfeasible. Alternatively, to characterize the non-equilibrium responsive behaviors of the system, one can introduce non-steady-state dynamics by increasing the demand level in a system that is in a steady state and measuring the relaxation time required for the system to establish a new steady state. In this study, the morning rush hour and evening rush hour (Supplement Meterials, B.3.1, Fig.B.16) are considered as two successive stimuli imposed on the system. Prior to the morning rush hour, the system is in a steady state where no stations with stranded passengers are observed. If the relaxation time of the system after the stimulus of morning rush hour is shorter than the time interval between the morning rush hour and evening rush hour, the system consistently returns to a steady state before the evening rush hour begins. This state is referred to as the normal state in present work. On the other hand, if the relaxation time is shorter than the time interval between the morning rush hour and evening rush hour, the accumulation of stranding phenomena within the system leads to a continuous decline in service efficiency. This state is referred to as the stranded state. By varying the passenger demand scale ($P_{r}$), the changes in average passenger waiting time and the 3 order parameters over time are plotted to observe the system’s responsive behaviors, as presented in Figure 4. The average passenger waiting time directly reflects the system’s service efficiency. To illustrate the phase transition characteristics, Figure 4A presents the summation of the average passenger waiting time over the entire day as a function of $P_{r}$. Importantly, a similar transition point, as shown in Figure 2A, is observed at $P_{r}\approx 1.4$, which can be defined as the threshold for resilience failure. When the demand scale is below this threshold, such as when $P_{r}=1.0$ representing the demand scale in a real normal day, 3 order parameters reflecting the level of stranded passengers exhibit distinct morning and evening rush hour characteristics with 2 peaks in the temporal dimension. Moreover, these 2 peaks do not overlap, creating a clear time window where no stranding phenomenon occurs. This finding implies that under normal demand scale for the morning rush hour and evening rush hour stimulus, the existing subway service effectively suppresses the occurrence of PPS within a limited time, without correlating to the subsequent rush hour. As the demand scale decreases, both the peak values of the 3 order parameters for the morning rush and evening hour diminish. When $P_{r}<1.0$, only the morning rush hour exhibits stations with stranded passengers and stranding clusters, while the evening rush hour is nearly absent. As $P_{r}$ increases, the relaxation time for the morning rush hour extends, and the start time of the evening rush hour is advanced. After $P_{r}$ reaches 1.4, the stranded station clusters generated by the morning rush hour fail to dissipate completely before the onset of the evening rush hour. Consequently, there is no time window without stranding between the morning rush hour and evening rush hour, making the subway network incapable of providing effective service due to PPS. This observation contrasts with the distinct dynamic characteristics of the morning rush and evening hour that are clearly separated when $P_{r}<1.4$. Hence, we define $P_{r}=1.4$ as the transition point of this system, serving as a metric for measuring threshold of service resilience failure.
Figure 4: (A) Average passenger waiting time ($t_{w}$) as a function of $P_{r}$. (B)-(D) Temporal evolution of max_cluster, n_stranded, and tot_stranded throughout the day as functions of $P_{r}$.
2.2.1 Dynamics Mechanism of Cluster Growth
Understanding the dynamic mechanism behind PPS generation is crucial for understanding the phase transition like behaviors in service systems and holds significant practical implications for enhancing the resilience of subway. The emergence of PPS in complex dynamical systems, comprising demand and supply basic units, is attributed to the complex and interconnected factors. In this study, the eigen microstate method of non-equilibrium statistical physics is used to uncover the characteristic dynamic patterns of PPS formation (see Method in the Appendix), which not only helps to understand the formation of clusters but also facilitates the study of cluster vanishing. include cluster formation and vanishing. According to the singular value decomposition (SVD)[39], the spatiotemporal correlation matrix A of stranded passenger number at each station can be factorized as $A=U\cdot\Sigma\cdot V^{T}$, where $U$ is the eigen microstate and $V$ is their evolution with time, $\Sigma$ is a diagonal matrix with the elements of $\sigma_{I}$. By equation ${{(A}_{I}^{e})}_{it}\equiv U_{I}\bigotimes V_{I}$, we get ${{(A}_{I}^{e})}_{it}$. We call the ensemble defined by ${{(A}_{I}^{e})}_{it}$ an eigen ensemble of system, where I represents the index of eigenstates, the subscript i denotes the station index, and t represents the time index. We consider $\sigma_{I}$ as the probability amplitude and $\omega_{I}=\sigma_{I}^{2}$ as the probability of the eigen ensemble ${{(A}_{I}^{e})}_{it}$ in the statistical ensemble A. The larger the probability $\omega_{I}$ associated with an eigen microstate, the greater its contribution to the formation of the PPS phenomenon. The eigen patterns with the largest eigenvalues are crucial factors determining the PPS phenomenon. By comparing the spatiotemporal characteristics of these key eigen microstates with the spatiotemporal characteristics of supply and demand, the micro mechanisms underlying the formation of these eigen patterns can be identified. Considering that the 4 four commuting corridors are independent in the spatial dimension, the eigen microstate method analyzing is applied separately to analyze Clusters I-VI.
Taking Cluster I (Fig.6A) of the northern commuting corridor as an example (Fig.6A), this corridor serves two adjacent large residential centers, Huilongguan and Tiantongyuan, collectively known as the "Huitian area" for commuting demands. This area covers approximately 63 square kilometers and accommodates nearly 900,000 residents, earning the title of "Asia’s largest sleeping town." A large number of residents commute between the Huitian area and the city center by taking subway as their daily transportation. We concentrate on the morning rush hour, from 6:00 to 13:00 to reveal the formation and vanishing of Cluster I. We take a time snapshot every 10 minutes to record the number of stranded passengers at each station in Cluster I, resulting in a spatiotemporal correlation matrix of stranded passenger counts with 43 time snapshots. By performing SVD on this matrix, we obtain 8 independent eigen microstates ${{(A}_{I}^{e})}_{it}$ and their eigen values $\omega_{I}$. 8 $\omega_{I}$ are shown in Fig. 6B , and the sum of probabilities of the top three micro eigenstates reaches 99.983%. We consider these 3 patterns as the major mechanisms of PPS phenomenon which are named based on their potential causes as: Local Pattern, Nonlocal Pattern, and Topology Pattern, respectively. Similar analyses are conducted for Cluster II-IV, as well, leading to evolution mechanisms similar to Cluster I (see Supplement Meterials, B.4.1, Fig.B.18,B.19,B.20). It is worth noting that the eigen microstate method can be regarded as an unsupervised machine learning method that excavates hidden patterns from the Spatiotemporal data. The interpretations of these patterns often require a phenomenological and empirical perspective[40]. Therefore, while the eigen microstate method used in this study provides rigorous eigen modes, the interpretation of the underlying micro mechanisms and the naming of these patterns are somewhat subjective and serve as a reference.
Figure 5: Station distribution and probabilities of eigen microstates for Cluster I. (A) Distribution of stations within Cluster I. The labeled stations include Station 200 (Huoying), Station 87 (LisuiQiao), Station 231 (BeiYuan), Station 232 (WangjingXi), Station 86 (Tiantong Yuan), Station 85 (Tiantong Yuan), Station 88 (Lishuiqiao Nan), and Station 89 (BeiyuanluBei). (B) Probabilities of the first eight eigen microstates corresponding to Cluster I during 6:00-13:00.
Local pattern: The largest eigenvalue, responsible for 84.88% of the total contribution, exhibits significant spatiotemporal variations at three distinct stations: i=87, 231, and 232. The temporal evolution of ${{(A}_{1}^{e})}_{it}$ at these stations is depicted in Figure 6A, demonstrating a characteristic pattern with a single peak during the early peak period, reaching its maximum around 8:50, and returning to the initial state by approximately 10:40. Station 231 exhibits the most prominent amplitude, while station 87 shows a more gradual variation. These observed patterns coincide with the temporal changes in the number of passengers entering the corresponding stations (i.e., demand scale), suggesting that this stranded pattern arises from an increase in the same passenger demand. Thus, we term this phenomenon the local pattern. Consequently, the primary driving factor behind the increase in stranded passengers at these stations is the mismatch between supply and demand caused by localized demand growth.
Figure 6: (A) Temporal evolution of the largest eigenstates in Cluster I for stations 87, 231, and 232 from 6:00 to 13:00. Stations 85, 86, 200, 88, and 89 do not exhibit significant variations in their eigenstates and are therefore not shown. (B) Variation in the number of entering passengers for stations 87, 231, and 232 from 6:00 to 13:00.
Nonlocal pattern: The second-largest eigenvalue accounts for 3.08% of the total variance. Figures 7(A), (C), and (E) depict the temporal evolution of ${{(A}_{2}^{e})}_{it}$ for stations 87, 231, and 232, respectively. In contrast to the characteristic single peak of Local pattern, this pattern exhibits alternating positive and negative variations, resembling the derivative of a single peak function. By comparing the morning rush hour stranded passenger counts at upstream stations i=86, 87, and 231 (Figures 7B, D, F) with ${{(A}_{2}^{e})}_{it}$, it can be inferred that the mechanism of stranded conditions involves propagation. This suggests that the variations in this pattern are driven by an increase in boarding passengers at upstream stations during rush hour periods. The stranded conditions at upstream stations result in a reduction of service supply downstream, leading to an increase in stranded passengers at downstream stations (Figure 1D). We named this phenomenon the Nonlocal pattern. Specifically, between 7:50 and 9:00, the increase in boarding passengers at station 86 leads to overcrowding, and this overcrowded condition does not improve at the nearest downstream station, 87. This interdependent relationship between adjacent stations facilitates the propagation of stranded conditions along the line direction. This effect is a secondary consequence of the Local pattern, resulting in a lower contribution from the Nonlocal pattern. Similarly, the increase in boarding passengers at station 231 exacerbates the stranded conditions downstream by reducing supply. During this period, both station 87 and station 231 experience stranded conditions, which have a suppressing effect on downstream station 232. Between 9:00 and 10:40, as the number of boarding passengers at station 231 increases, the stranded conditions at downstream station 232 worsen. Simultaneously, the decrease in boarding passengers at station 86 gradually reduces the occupation of supply at downstream stations, thereby alleviating the stranded conditions. Additionally, the impact of station 87 on the stranded conditions at station 231 starts to weaken. This mode reflects the propagation of stranded conditions and explains how the traffic supply at upstream stations influences the stranded conditions at downstream stations.
Figure 7: Variation of the second-largest eigenstates and passenger demand in Cluster I. (A), (C), and (E) depict the temporal evolution of the second-largest eigenstates for stations 87, 231, and 232, respectively, from 6:00 to 13:00. (B), (D), and (F) represent the changes in the number of boarding passengers at upstream stations 86, 87, and 231, respectively. The eigenstates of other stations within Cluster I exhibit minimal variations.
Topological pattern: The third-largest eigenvalue corresponds to a pattern contribution of 1.99%. Fig. 5 illustrates the temporal evolution of ${{(A}_{3}^{e})}_{it}$. This pattern unveils the influence of the infrastructure’s topological structure on passenger stranding, which we refer to as the Topological pattern. The distinguishing characteristic of this pattern is the significant variation observed at a specific station, designated as Local pattern, as depicted in Fig. 9 for station 87. Unlike the other 2 stations, station 87 functions as a transfer station, with multiple upstream stations contributing to the propagation of stranding and multiple downstream stations contributing to stranding alleviation, while stations 231 and 232 serve as non-transfer stations. Consistent with the findings from steady-state condition calculations under spatially and temporally homogeneous demand conditions discussed in the previous section, stations with high network betweenness centrality at transfer points play a pivotal role in PPS occurring. Under steady-state conditions, only the topological structure of the network influences PPS. In this system characterized by spatially and temporally heterogeneous supply and demand, the contribution of the topological factor ranks third.
Figure 8: Temporal evolution of the third-largest eigenstates for stations 87, 231, and 232 from 6:00 to 13:00.
2.2.2 Validating the Mechanism of Cluster Formation
The match between supply scale and demand scale is a fundamental factor determining the state of a system. To validate the interpretation of 3 pattern of cluster formation, addition simulations with varying the supply scale and fixing the demand scale are performed. The supply scale is controlled by varying the train frequencies on both Line 5 and Line 13, where Cluster I is located, while maintaining the spatiotemporal distribution characteristics of the supply. This allowed us to observe the inhibiting effect of increased supply scale on the both Local and Nonlocal patterns. Detailed information on the train departure frequency control scheme can be found in Supplement Meterials, B.4.2. By employing the eigen microstate method, the Local, Nonlocal, and Topological patterns of Cluster I under different service supply scales Ps (see Supplement Meterials, B.4.3, Fig.B.22) are analyzed. As shown in Fig. 9A, an increase in the supply scale leads to a continuous rise in the contribution of the Topological pattern, while the contributions of the Local and Nonlocal patterns are suppressed. In these simulations, the topological structure of the network remains unchanged, and any absolute changes in stranded passengers resulting from the topological structure can be neglected. The increase in the contribution of the topological pattern in Fig. 10A is only attributed to the decreased contributions of the local and nonlocal patterns. This supports our reasonable assumption that the local and nonlocal patterns arise due to a mismatch between supply scale and demand scale.
When $P_{s}>1.5$, the Nonlocal pattern, an important mechanism in PPS, is almost completely suppressed, and the contribution of the Local pattern is significantly reduced. Compared to the Nonlocal pattern, the local pattern is more sensitive to changes in supply scale. When $Ps\approx 1.3$, the contribution of the local pattern rapidly decreases, indicating that the increased supply scale directly fulfills the local supply shortage. During this process, the contribution of the Nonlocal pattern decreases slowly. The Local pattern is highly sensitive to supply improvements, hence an increase in demand leads to a decrease in the Local pattern. The triggering factor for the Nonlocal pattern is the presence of local stranding, which allows the propagation of the Nonlocal pattern through the PPS mechanism to downstream stations. Therefore, the regulation of the Nonlocal pattern by the supply scale is indirect, with the response to supply increases occurring later than that of the Local pattern. It is worth noting that when the supply scale exceeds 1.5, the contribution of the Nonlocal pattern is almost completely eliminated, and further improvements in the supply scale lead to slower enhancements in the Local pattern.
The results validate the reasonableness of the cluster formation mechanism analyzed above and would provide guidance for decision making of policies. Increasing the supply scale within a certain range is beneficial for suppressing both the Local and Nonlocal patterns, and the influence on the Nonlocal pattern gradually strengthens as the supply increases between 1.3 and 1.5. After the supply scale exceeds 1.5, the inhibiting effect of the Local pattern significantly weakens, and the Nonlocal pattern is completely eliminated. Increasing the supply scale has limited effect on suppressing the Local and Nonlocal patterns. If further improvement in the system’s service efficiency is desired, it is necessary to shift the regulatory policies towards optimizing the network’s topological structure.
Figure 9: The proportion of local pattern (green), nonlocal pattern (red), and network topology pattern (blue) in the total variance as a function of $P_{s}$.
3 Conclusion
In this study, we examine the subway system as a complex system that consists of 2 type of basic units, namely passengers and subway trains with a service relationship. By employing computational simulations of a big-data-driven model system, we explore the dynamics of stranding phenomenon within the subway system, particularly how it evolves in response to variations in the scale of supply. Our computational results unveil analogous behaviors of phase transitions in both service efficiency, as quantified by the average waiting time for passengers, and the scale of stranding phenomenon within the subway. We identify a transition point, which quantitatively characterizes the threshold of subway service resilience failure. Employing the method of eigen microstates from non-equilibrium statistical physics, we elucidate the formation mechanisms of PPS phenomenon. These achievements enhance the understanding of the features of service efficiency in subway networks and provides a theoretical foundation for strategies such as topological planning and optimization of train schedules to improve service. The proposed framework, involving complex systems comprising multiple interdependent basic units with service relationships, offers a valuable perspective for studying urban systems. In the pursuit of smart cities, this framework serves as a reference for evaluating urban planning, governmental policies, and business decisions. Furthermore, the methodology used in present work has broader implications for similar scheduling and logistics management challenges encountered in other service systems, including demand-driven or flexible bus scheduling, elevator dispatching in skyscrapers, supply chain management, medical appointment assignment, logistics management for aircraft carriers, and container logistics management.
Supplement Meterials
Appendix A Data and System
A.1 Data Description and Processing
A.1.1 Data Description
(1) Automatic Fare Collection data In this study, the actual passenger demand dataset for the service system is obtained from the Automatic Fare Collection (AFC) system of Beijing Rail Transit. The dataset comprises passenger travel data collected by the AFC system and includes all travel information from May 10th to May 14th, 2021. A total of 6,734,988 passenger records are obtained. The data attributes include TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME, PRODUCT ISSUER ID, PRODUCT TYPE, PAYMENT VALUE, CARD LIFE CYCLE COUNT, RECONCILIATION DATE, SETTLEMENT DATE, SAM ID, DEVICE ID, SOURCE PARTICIPANT ID, PURSE REMAINING VALUE, PURSE_REMAINING_VALUE, PTSN. The data used in this study are TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME.
(2) Subway Train Timetable Data
      (a)Train Classifications: The train in Beijing are categorized as A, B, C, D, and L. In the network model, all train on the lines are of type B, with a maximum capacity of 250 passengers per carriage. In the deal system: The passenger capacity of each train in the model is set at 1500 passengers per train. In the realistic model: The train models on each line in the network model are identical, but the train compositions are not entirely consistent on different lines. The maximum passenger capacity is solely determined by the train model and composition, resulting in slight variations in the maximum capacity among different lines. The train carriage models, compositions, and maximum passenger capacities for each line are listed in Table A.1. The actual capacity of the train carriages in the simulation system needs to be adjusted through parameter tuning, and the tuning results are presented in the following sections.
Table A.1: Train Type, Composition, and Maximum Passenger Capacity Information
Line Composition Maximum passenger capacity
1 6 1500
2 6 1500
4 6 1500
5 6 1500
6 8 2015
7 8 2015
8 6 1500
10 6 1500
13 6 1500
(b) Categorized by Travel Segments: Shuttle trains and non-shuttle trains. The realistic system only includes shuttle trains, while the ideal system encompasses both shuttle and non-shuttle trains.
(3) Train Timetable Data The train timetable data represents the actual data for trains on the respective day, with the number of train services varying across different lines as shown in Table A.2. The Beijing subway timetable data used in this study primarily comes from the official application of Beijing Rail Transit App. This official app, maintained and released by the Beijing subway operating authority, provides comprehensive subway timetable information. The timetables include essential information such as train departure and arrival times, station details, and train frequencies for each line of the Beijing subway system.
Table A.2: The total number of train services on different lines on May 10th, 2021
Line
1
2
4
5
6
7
8
10
13
total number of train services
652
543
667
594
614
760
394
495
574
A.1.2 Data Processing
(1) AFC Data
      In this study, the network model consists of 9 out of all the lines in Beijing subway network. To process the original dataset, the following steps were taken:
      (a) Removal of records with missing key information, such as missing entry time, exit time, entry location, or exit location.
      (b) Elimination of duplicate records for the same station at the same time, both for entry and exit.
      (c) Exclusion of records with identical passenger ID, entry swipe time, exit swipe time, entry location, and exit location.
      (d) Exclusion of transaction records generated before 4:30 AM or after 12:30 AM, corresponding to the operational hours of the urban rail transit system.
      (e) Selection of records with entry and exit locations limited to stations on Lines 1, 2, 4, 5, 6, 7, 8, 10, and 13.
After these processing steps, a total of 3,811,418 AFC data transaction records that meet the research requirements were obtained. (2) Train Travel Data The train timetable data used in this study was obtained through web scraping of real-time data from the Beijing Rail Transit App. The data covers the period from May 10th to May 14th, 2021. Based on the acquired train timetable data, the departure time for each train was extracted. Using the timing information for the arrival at two consecutive stations, the travel time between each pair of stations was calculated as the time interval. These calculated intervals represent the travel time between stations.
A.2 2 Types of Systems
We have constructed an ideal system and a realistic system, with each system consisting of three components: the operation of trains, the generation of passengers, and the trains’ service to passengers.
A.2.1 The operation of trains
(1) Departure
Once the departure time is reached, the system automatically generates new trains from each originating station, without considering the source of the trains. The departure time, originating station, and destination station are set as follows:
      (a) In the ideal system The originating and destination stations for the trains are obtained from actual data (Table A.3), with each line having two directions of travel. On linear lines, the originating and destination stations are located at the two ends of each line. On circular lines, the originating and destination stations are adjacent stations. The direction of departure is from the originating station towards the destination station. From the start of the simulation, trains on each line start departing from their respective originating stations and continue until the end of the simulation, with a departure interval set at 4 min.
Table A.3: Information of Line Types, Direction, Originating Stations, and Destination Stations
Line Type Direction Type Originating Stations Destination Stations
1 Straight 1 Gucheng Universal Report
2 Universal Report Gucheng
2 Circle 1 Jishuitan Xizhimen
2 Xizhimen Jishuitu
4 Straight 1 Anheqiaobei Tiangongyuan
2 Tiangongyuan Anheqiaobei
5 Straight 1 Tiantongyuanbei Songjiazhuang
2 Songjiazhuang Tiangongyuanbei
6 Straight 1 Tiantongyuanbei Songjiazhuang
2 Songjiazhuang Tiangongyuanbei
7 Straight 1 Liuliqiaodong Universal Resort
2 Universal Resort Liuliqiaodong
8 Straight 1 Yinghai Zhuxingzhuang
2 Zhuxingzhuang Yinghai
10 Circle 1 Bagou Huoqiying
2 Huoqiying Bagou
13 Straight 1 Xizhimen Dongzhimen
2 Dongzhimen Xizhimen
(b) In the realistic system
      The departure time, originating stations, and destination stations for both shuttle and non-shuttle trains on each line are obtained from actual data.
(2) Travel Process
      Trains always travel on a single line, starting from the originating station and sequentially passing through each station between the originating station and the destination station until reaching the destination station.
      Trains travel between adjacent stations, maintaining a constant speed during the travel process. The travel time is determined by the actual distance between stations. The actual interval data between stations is set as follows. The departure and arrival times of trains at each station in both directions are obtained from actual data. The line segments between stations are abstracted as edges between nodes, and following the directions specified in Fig. A.10, each edge on each line is numbered sequentially as "1, 2, 3, …" until all segments on the line are numbered. Based on the timetable data, the time intervals ($T_{i}$) between all stations in both directions of each line are calculated. Only the intervals satisfying $Ti\in(0,1000]$ seconds are selected. This time interval includes two components: the travel time of trains on the edges ($T_{r}$,0) and the service time to passengers ($T_{s}$) (set as 30s). By subtracting the service time ($T_{s}$) from the total interval, the travel time ($T_{i,1}$ and $T_{r,2}$) of trains on each edge of the line in both directions is obtained. The average of travel times on corresponding edges of both directions is calculated to obtain the travel time ($T_{i}$) of trains on each edge in the simulation system (Fig. A.10). This process starts from the departure and continues until all stations on the line have been traversed. After completion, trains are removed from the system, and post-travel scheduling is not considered.
Figure A.10: Travel time between stations. The direction of lines in the table is arranged according to the direction from originating station to destination station (direction 1) on each line.
A.2.2 The Generation of Passengers
(1) Setting of Passenger Generation Time, Originating Stations, and Destination Stations
      The data input for the system is as follows: the passenger’s entry time serves as the time when the passenger is generated in the simulation, while the entry station and exit station represent the originating and destination stations in the system, respectively. The following describes the simulation of passengers in both the model system and the realistic model.
      (a) In the ideal model:
      At the start of the simulation, the system generates $n_{p}$ passengers every 1 min. The originating stations and destination stations for the generated passengers are randomly selected independently from $N$ stations ($N=234$). The passenger demand in the model system is denoted as $P_{u}$ and calculated as $n_{p}/N$.
      (b) In the realistic model:
      For the realistic passenger demand, the entry time and passenger OD are obtained from actual data, and the corresponding demand scale is set to 1.0.
      For additional passenger demands, $N_{d}$ instances are randomly selected from the original $N_{r}$ passenger data (each instance containing entry time, originating station, and destination station) to represent the modified entry passenger information, resulting in $N_{d}$ instances of passenger data. The adjusted passenger demand, denoted as $P_{r}$, is calculated as $N_{d}/N_{r}$.
      When the simulation time reaches the entry time of a passenger, the system generates the corresponding passenger at the originating station, and the passenger immediately enters the waiting state.
(2) Passenger Route Assignment
      We employ the Breadth-First Search (BFS) algorithm to find $k$ ($k=5$) shortest paths for each origin-destination (OD) pair. These paths are evaluated based on three factors: the number of stations passed ($n_{s}$), the total travel time ($T=T_{d}+T_{l}$, where $T_{d}$ represents the total travel time and $T_{l}$ represents the total station dwell time), and the number of transfer stations ($n_{t}$). A scoring mechanism is applied to the k paths, with the highest-scoring path selected as the route for that OD pair. The scoring rule is defined as follows:
$Score=(10/n_{s}+1000/T+1/n_{t})/111.$ (A.1)
A.2.3 Trains’ Service to Passengers
(1) Service Process and Rules       Prior to the simulation, the number of trains, timetables for each train, and maximum passenger capacity per train have been remain unchanged throughout the simulation. For each arriving train, the following rules are followed in a specific order: first, passengers disembark before boarding; priority is given to passengers who arrived first, and if a train is already at maximum capacity, passengers cannot board. Specifically, the operations for trains stopped at stations are as follows:
      (a) The process of passengers getting off
      If the train reaches its final destination, all passengers onboard get off. Among the alighting passengers, those whose alighting station differs from their intended destination station need to transfer, otherwise they are removed. If the train has not reached its final destination, passengers who have reached their intended destination station need to get off and are removed. Among the remaining passengers who have not reached their destination, those whose next station on their path differs from the next station the train will arrive at need to transfer. The transfer rule is as follows: after alighting, passengers who need to transfer must queue for their next train on the platform that aligns with both the passenger and the train’s next station.
      (b) Passenger Boarding
      If a passenger’s next destination matches the train’s next station, they board the train in a first-come, first-served order. Otherwise, the passenger continues to wait in the queue.
      (c) Simulation Duration
      The actual system duration is from 0 to 50400s and is collected every 5min.
      For the realistic system covered the period from 4:30 AM to 12:30 AM (a total of 73800s), data is collected every 10 min. (2) Dynamic Process of Passenger Boarding and Disembarkation The dynamic process of passenger boarding and alighting in the system can be illustrated by Fig.A.11, with specific symbols explained in Table A.4.
      We introduce $P^{\alpha}$ to represent the state of the train, and $P_{b}$ represents the station where the train is located. $P^{\alpha}=0$ indicates that the train has not yet departure, $P^{\alpha}=-1$ indicates that the train is loading passengers at the station, $P^{\alpha}=1$ represents the train is in travelling, and $P^{\alpha}=2$ indicates that the train has reached its destination. Passengers in the system are generated as follows. A passenger is generated every 1min, and their origin and destination are randomly selected from all stations. The number of generated passengers is set based on actual demand. Prior to the simulation, the passenger’s entry time, origin, and destination are set. During the simulation, the system utilizes the K-shortest path algorithm to determine the best travel route for passengers based on their origin and destination at the time of entry. The system records the travel demand status $S^{\alpha}$ for all generated passengers at each moment. There are four specific states: $S^{\alpha}=-1$ represents passengers waiting for boarding at the platform, $S^{\alpha}>0$ indicates that passengers have boarded and $S^{\alpha}$ represents the train number they are on, and $S^{\alpha}=0$ represents passengers have reached their destination.
Figure A.11: Passenger Boarding and Alighting Dynamics.
Table A.4: Notations, meaning and comments of the system
Notation Meaning Comments
$P^{\alpha}$ the state of trains $P^{\alpha}=-1$: loading passengers; $P^{\alpha}=1$: traveling; $P^{\alpha}=2$: arriving at destination.
$S^{\alpha}$ the state of passengers $S^{\alpha}=-1$: waiting; $S^{\alpha}>0$: being onboard, and $S^{\alpha}$ is the number of the train, $S^{\alpha}=0$: arriving at the destination.
$P^{s}$ the state of stations $P^{s}=1$: arriving at the destination; $P^{s}=2$: arriving at transfer stations, $P^{s}=0$: otherwise.
$P^{\beta}$ the location of the train current station number of the train
$P^{n}$ next station of the train next station number of the train
$P^{c}$ occupancy status of the train $P^{c}=1$: fully loaded; $P^{c}=0$: not fully loaded
$S^{d}$ passengers’ destination Passengers’ destination
$S^{n}$ next station of passengers next station of passengers
A.2.4 Parameter settings and validation
Due to the fact that the network model does not include all the routes in Beijing’s entire transportation network, there are passengers in the real traffic flow whose origins or destinations are not located on the 9 lines of the network model. This leads to some deviations between the simulated system’s train capacity and the actual train capacity. To address this issue, we utilize the entry and exit timestamp from the AFC data to determine the travel time for passengers. By comparing the simulated travel time with the actual travel time derived from the AFC data, we adjust the parameter of train car capacity.
      Since different types of trains operate on the actual routes, there are differences in train capacities among different lines. The train capacities for each line are listed in Table A.1. During the parameter adjustment process, we maintain the proportional relationship between the car capacities of different lines consistent with the actual train capacities on each line. Let q0 represent the reference capacity of trains for Line 1, and q represents the adjusted train car capacity for Line 1, as shown in Table A.5.
Table A.5: Information on Reference Capacity and Design Standard Passenger Capacity of Subway Trains on Different Lines
Line Reference Capacity Standard Passenger Capacity
1 1500 1460
2 1500 1460
4 1500 1460
5 1500 1460
6 1500 1960
7 2014 1960
8 2014 1460
10 2014 1460
13 2014 1460
We calculated the fitting effect of different train capacity models by following steps.
      Step 1: We obtained the actual boarding time $t_{s,r}$ and alighting time $t_{e,r}$ for each passenger from the AFC data. We also obtained the corresponding passenger boarding time $t_{s,s}$($t_{s,r}=t_{s,s}$) and alighting time $t_{e,s}$ from the simulation system.
      Step 2: The actual travel time for each passenger is given by $\Delta t_{r}=t_{e,r}-t_{s,r}$, and the simulated travel time is given by $\Delta t_{s}=t_{e,s}-t_{s,s}$. The absolute error between the simulated and actual travel times is calculated as $\Delta t=\Delta t_{r}-\Delta t_{s}$.       Step 3: We applied the 3-sigma criterion to remove outliers from the set of absolute errors.
      Step 4: The absolute time errors are normalized to obtain the relative error for each passenger, defined as $\Delta t^{\prime}=\Delta t/\Delta t_{r}$. A smaller value of $\Delta t^{\prime}$ indicates that the simulation results for that passenger are closer to the actual data.
      Step 5: We calculated the mean $M_{t}$ and standard error ${SD}_{t}$ of the set of relative errors $\Delta t^{\prime}$. A mean value and variance closer to 0 indicate that the simulation results are closer to the actual data.
      By comparing the mean and variance of the relative errors for different train capacities (Fig. A.12), we observed that as the car capacity increases, both $M_{t}$ and $SD$ decrease. When $q=1.0$, $M_{t}$ and $S_{t}$ approach 0, indicating that the system corresponding to this set of car capacities is very close to the real system.
      In addition, we consulted the "Urban Rail Transit Engineering Project Construction Standard" (JB 104 2008), which specifies that in the case of overcrowding, the seating capacity remains unchanged, while the standing capacity is calculated as 9 persons/m2. Type B trains are comprised of 2 types of trains: single-driver trains and non-driver trains. The overcrowding threshold for single-driver trains is set at 230, while the overcrowding threshold for non-driver trains is set at 250. Each train set includes a single-driver train at each end, with the remaining cars being non-driver trains. The studied subway lines 1, 2, 4, 5, 6, 7, 8, 10, and 13 all use B-type trains, with the exception of lines 6 and 7, which have 8-car sets, while the rest of the lines have 6-car sets. Based on this information, we obtained the maximum passenger capacity for each line and train set (see Table A.5), which closely matches the parameter setting in our model, validating the effectiveness of our parameter adjustments.
Figure A.12: Variation of the Mean $M$ and Variance $SD$ of the Relative Error Set $\Delta t^{\prime}$ for the Simulated System with Different Train Capacities $q_{c}$. The blue curve represents the change in the mean $M$, and the orange curve represents the change in the variance $SD$. A closer approach to 0 for both $M$ and $SD$ indicates that the simulation model’s performance is closer to the actual data.
Appendix B Supplementary text
B.1 Quantitative definition of cluster
The number of stranded passengers at station $i$ at time $t$ is denoted as $S_{i}(t)$. The relation between two stranded stations, $i$ and $j$, can be classified into two categories:
$e_{ij}=\begin{cases}1&\text{ $S_{i}S_{j}>0$ and $T_{ij}=1$ }\\ 0&\text{ else }\end{cases}$ (B.2)
The spatial correlation between station $i$ and station $j$ along the route is denoted by $T_{ij}$: $T_{ij}=1$ if there is a station, $S_{i}\neq 0$, that establishes connectivity between stations $i$ and $j$ on the route; otherwise, $T_{ij}=0$. The propagation of stranding in the spatial domain and the formation of clusters composed of stranded stations are determined based on $e_{ij}$.
B.2 System I: Steady-State features
B.2.1 Time Translation Invariance of the Model
When the demand scale is relatively small, such as $P_{u}=20$, the evolution of the three order parameters over time is shown in Fig. B.13a. In the initial stages of the simulation, all vehicles in the system depart from the originating station, while passenger demands are randomly generated throughout the network. Due to insufficient service supply in the early stages, all three order parameters increase rapidly. However, after a relaxation time of $2.0\times 10^{4}$, they reach a steady state. As the demand scale $P_{u}$ increases (as shown in Fig. B.13B for $P_{u}=30$), the system reaches a steady state as long as the relaxation time is sufficiently long. In the steady state, all the order parameters significantly increase with the increasing demand scale $P_{u}$. We use the average values of the 3 order parameters after the system reaches a steady state to characterize the state of the system under a given passenger demand scale $P_{u}$ condition.
Figure B.13: (A) The variation curves of n_congestion, max_cluster, and tot_congestion in the system when $P_{u}=20$. (B) The variation curves of n_congestion, max_cluster, and tot_congestion in the system when $P_{u}=30$. In fact, for all $P_{u}$ values in the system, the four variables eventually reach a steady state. For clarity, only the changes in the number of clusters in the system for two $P_{u}$ values are plotted.
B.2.2 Distribution Evolution of Station with stranded passengers in ideal model
Figure B.14: Evolution of station distribution for the ideal model with $P_{u}=20$.
Figure B.15: Evolution of station distribution for the ideal model with $P_{u}=30$.
B.3 System II: Spatiotemporal Mismatch between Supply and Demand
B.3.1 Tidal characteristics of demand and supply
Figure B.16: (A) Temporal distribution of demand and supply. Note: $N_{supply}$ represents the total number of trains departing every 10 min in the model shown in Fig. A.11, and $N_{need}$ represents the total number of passengers entering the station every 10 min in the model shown in Fig. A.11. (B) Spatial distribution of passenger demand between 6:00 and 13:00. The diameter of the circles is proportional to the number of passengers entering the station at each location. Larger circles indicate higher passenger volumes, while smaller circles indicate lower volumes.
B.3.2 Analysis of Cluster Evolution
In the analysis of microstate, we employ the following statistical methods for the variables in the system.
(1) Number of stranded passengers: We introduce the variable $S_{i}^{t}$ to record the stranded status of passenger i. If a passenger fails to board the subsequent train after entering the station, they are considered stranded, and $S_{i}^{t}=1$; otherwise, $S_{i}^{t}=0$. We calculate the total number of stranded passengers at each station as the current station’s count of stranded individuals. This count is updated every 10 minutes.
(2) Number of passengers entering each station: This data is obtained from the raw data, and we calculate the number of passengers entering each station at every time step.
(3) Number of passengers boarding: Since train arrivals at stations are discrete events, we adopt a method to account for variations caused by the presence or absence of trains during the data collection. If a train is currently in the passenger-loading state at a particular station, the number of boarding passengers is determined by summing the passenger count from that train’s arrival time up to the current time. Otherwise, it is determined by the number of passengers carried by the last departing train from that station.
B.3.3 Clusters and Labeling
During the evolution process of the real supply-demand system, if adjacent stations have experienced stranded passengers, they are considered to belong to the same cluster. During the morning peak hours on weekdays, four stranded station clusters have formed in four directions in Beijing (see Fig. B.18). The stations included in each cluster are listed in Table B.6 and Table B.7.
Figure B.17: (A)-(D) Spatial distribution of stranded stations at 7:20, 8:00, 8:20, and 13:00 corresponding to $P_{r}=1.0$. (F)-(I) Spatial distribution of stranded stations at 6:40, 7:30, 8:20, and 13:00 in the network corresponding to $P_{r}=1.6$. (E) and (J) Clusters formed by stations that experienced stranded situations during the morning peak hours (6:30-13:00). These clusters consist of adjacent stations in both spatial and temporal dimensions. The stations included in Cluster I-IV 4 are represented by green, blue, yellow, and purple colors, respectively.
Table B.6: Stations and Station Labels included in the 4 Stranded Station Clusters Formed during Weekday Morning Peak Hours with Passenger Demand of 1.0.
Cluster[c] Station(Label)[c]
Cluster1 Huilongguan (230), Yuxin(200), Lishuiqiao(87), Beiyuan(231), Wangjingxi(232)
Cluster2 Jintailu(119), Shilipu(120), Qingnianlu(121)
Cluster3 Hepingmen(46), Xuanwumen(47), Caishikou(67), Taoranting(68), Beijing South Railway Station(69), Majiapu(70), Jiaomenxi(71), Gongyi Xiqiao(72), Xingong(73), Xihongmen(74), Gaomidianbei(75), Gaomidiannan(76)
Cluster4 Fuxingmen(1), Fuchengmen(4), Chegongzhuang(50), Chegongzhuangxi(114),Baishiqiaonan(11), Huayuanqiao(112), Cishousi(111), Haidian Wuluju(110)
Table B.7: Stations and Station labels included in the 4 Clusters formed during weekday Morning rush hour with $P_{r}=1.6$.
Cluster[c] Station(Label)[c]
Cluster1 Huilongguan(230), Yuxin(200), LishuiQiao(87), Beiyuan(231), Wangjingxi(232), Huoying(119), Tiantongyuannan(86)
Cluster2 Jintailu(119), Shilipu(120), Qingnianlu(121), Dalianpo(122), Huangqu(123), Changying(124), Yong’anli(18), Guomao(19), Dawanglu(20), Sihui(21), Sihuidong(22), Gaobeidian(23), Communication University of China(24), Shuangqiao(25), Guanzhuang(26), Baliqiao(27), Tongzhou Beiyuan(28), Guoyuan(29)
Cluster3 Hepingmen(46), Xuanwumen(47), Caishikou(67), Taoranting(68), Beijing South Railway Station(69), Majiapu(70), Jiaomenxi(71), Gongyi Xiqiao(72), Xingong(73), Xihongmen(74), Gaomidianbei(75), Gaomidiannan(76), Caoqiao(159), Zaoyuan(77), Qingyuanlu(78)
Cluster4 Fuxingmen(11), Fuchengmen(49), Chegongzhuang(50), Chegongzhuangxi(114), Baishiqiaonan(113), Huayuanqiao(112), Cishousi(111), Haidian Wuluju(110), Xizhimen(51), Fuxingmen(10), Muxidi(9), Military Museum(8)
B.4 Validating the Mechanism of Cluster Formation
B.4.1 Validating the Mechanism of Cluster II-IV
Figure B.18: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster II.
Figure B.19: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster III.
Figure B.20: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster IV.
B.4.2 Train Control Scheme
We maintain the spatial distribution of train departures while adjusting the frequency of train departures on Lines 5 and 13 to modify the supply scale. Initially, we preserve the spatial distribution of train departures according to the original timetable for Lines 5 and 13. We double the train departure density, resulting in a total number of train services denoted as $N_{tot}$. The additional trains in the expanded timetable are scheduled to depart 2 min later than the original timetable (Fig. B.21). Since the time delay is relatively short, the new timetable preserves the spatiotemporal characteristics of the original one. Based on the new train timetable, we randomly select the required number of train services according to passenger demand to form the actual subway timetable. Let $N_{ori}$ denote the total number of train services in the original timetable. We define the adjusted supply, $P_{s}$, as $N_{tot}/N_{ori}$.
Figure B.21: Changes in the Number of Train Services for the Original Timetable ($N_{o}ri$), Additional Timetable ($N_{a}dd$), and New Timetable ($N_{t}ot$) for Lines 5 and 13.
B.4.3 Changes in the Contributions of Different Supply Scale Patterns
Figure B.22: Proportional contributions of Local (green), Nonlocal (red), and Topological (blue) Patterns for different $P_{s}$.